Given that 𝑓(βˆ’1)=3and 𝑓′(βˆ’1)=2,find an equation for the tangent line to the graph of 𝑦=𝑓(π‘₯)at π‘₯=βˆ’1

.



1
Expert's answer
2021-05-07T09:04:11-0400

We know that the derivative indicates the slope of the tangent line:


"m=f'(-1)=2"

Therefore, the tangent line requested is the line that passes through the point "(-1,3)" and has "m=2" as slope:


"y-3=2(x+1)\Rightarrow y=2x+5"